OpenStudy (lovelycharm):

The tip of a 12-inch wiper blade wipes a path that is 30 inches long. What is the angle of rotation of the blade in radians to the nearest tenth?

OpenStudy (retireed):

Calculate the circumference of the circle. Calculate the ratio of the arc sweep of the circle in radians.

OpenStudy (lovelycharm):

so you calculate both number

OpenStudy (retireed):

The circumference is C = 2 (pi) r or 24 pi

OpenStudy (retireed):

The arc sweep is 30 inches

OpenStudy (retireed):

The ratio is 30 inches / 24 pi = 30 / 75.398 = 0.3979ths of the circle

OpenStudy (lovelycharm):

so you round of right

OpenStudy (retireed):

I round just for typing, I usually carry it all on the calculator.... Convert 0.3979 of a circle to radians

OpenStudy (lovelycharm):

Okay thank you so it would be 0.4

OpenStudy (retireed):

No you need to express 0.3979 of a circle in radians

OpenStudy (lovelycharm):

i got it 1.3

OpenStudy (retireed):

A circle is 2 pi radians so 0.3978 * 2 is 0.796 pi I don't like working in radians since I have never been very good at it.

OpenStudy (retireed):

I got you as close as I can figure out, sorry.

OpenStudy (lovelycharm):

it okay thank you